OpenStudy (anonymous):

what is 2x-3(x+4)= -11

sam (.sam.):

2x-3(x+4)=-11 2x-3x-12=-11 -x-12=-11 -x=1 x=-1

OpenStudy (anonymous):

okay. thank you. i appriciate it